OpenStudy (dls):

Venus has a mass of about 0.558 times the mass of earth and a diameter of about 0.381 times the diameter of earth .The acceleration of a body falling near the surface of venus is...?

OpenStudy (anonymous):

\[g=G\frac{M}{r^2}\\ \frac{g_v}{g_e}=\frac{M_v/r_v^2}{M_e/r_e^2} \]

OpenStudy (dls):

r=0.19?

OpenStudy (anonymous):

rv/re=0.19

OpenStudy (dls):

i used the same formula,but mgetting the wrong numerical value

OpenStudy (anonymous):

Did you forget to square the r? Show your work if possible.

OpenStudy (dls):

i did..let mewrite

OpenStudy (dls):

Gv=GMe(0.558)/(Re0.19)^2 Ge=GMM/Re^2 => 9.8*0.06/0.2*0.2 (approxed)